OpenStudy (anonymous):

Are the graphs of 3x-4y=16 and y=3/4(x-1) parallel? Explain.

OpenStudy (anonymous):

Compute the slopes of both lines. If they're the same, the lines are parallel.

OpenStudy (anonymous):

They are parallel. You just need to be able to recognize that their slopes are the same. It would probably help if they were all in the same form, such as y=mx+b.

OpenStudy (anonymous):

3x-4y=16 and y=3/4(x-1) are parallel

OpenStudy (anonymous):

3x-4y=16 dividing by 4 we get (3/4)x-4y=4 (3/4)x-4=4y dividing by 4 (3/16)x-1=y here slope m=3/4 and y=3/4(x-1)